An iron rod is 1.58m long at 0oc. What must be the length if it is heated through temperature of 60oc [iron linear expansivity = 0.000012 k]
1
Expert's answer
2020-04-26T18:51:02-0400

The linear thermal expansion equation says

"l(t)=l(t_0)(1+\alpha(t-t_0))\\
l(60^{\circ}\rm C)=1.5800(1+0.000012\times 60)=1.5811\:\rm m."